An electric lamp of mass 900g ,hungs on a flex which is pulled aside by a horizontal string until a flex makes angle of 30 With vertical. Find tension in the string and in a flex
1
Expert's answer
2021-09-04T15:19:29-0400

Solution.

"m=0.9kg;"

"\alpha=30^o;"

"T_f=\dfrac{mg}{cos\alpha};"


"T_f=\dfrac{0.9\sdot9.8}{0.87}=10.14N;"

"T_s=mgtan\alpha" ;

"T_s=0.9\sdot9.8\sdot0.59=5.2N;"

Answer: "T_f=10.1N; T_s=5.2N."